OpenStudy (anonymous):

if f(x)=x2-7x find f(a-4)

hero (hero):

f(x) = x^2 - 7x f(a-4) = (a-4)^2 - 7(a -4) f(a - 4) = (a - 4)(a - 4) - 7(a - 4) f(a -4) = a(a - 4) - 4(a - 4) - 7(a - 4) f(a - 4) = a^2 - 4a - 4a + 16 - 7a + 28 f(a - 4) = a^2 - 4a - 4a - 7a + 16 + 28 f(a - 4) = a^2 -(4 + 4 + 7)a + 44 f(a - 4) = a^2 -15a + 44

OpenStudy (anonymous):

my question is did you take a-4 to replace the x in each one

hero (hero):

Yup. I'd be more impressed if you understood the steps as well.

OpenStudy (anonymous):

i know it also be nice if the school i am at would match the problems up with the chapters they give us each week